OpenStudy (anonymous):

Plesae help me....... I have the answer can somone please check me.... Find values for a and b so that f(x)=x^3+ax+b has a relative minimum at (2,-9) a=? b=?

OpenStudy (anonymous):

a=-12 b=4

OpenStudy (campbell_st):

so the point (2, -9) needs to be on the curve so if you substitute into the equation you get -9 = (2)^3 + 2a + b or -17 = 2a + b you also knoq that (2, -9) is a minimum... so the value of the gradient at the point is zero. so the 1st derivative is f'(x) = 3x^2 + a and if f'(x) = 0 when x = 2 then 0 = 3(2)^2 + a or - 12 = a substitute it into the original equation to find b.